Krissy525:

I need some help with two of my math problems. The side length of a cube can be represented by the expression 2x^5. If the side length is doubled, write an expression to represent the new volume of the cube. The area of a triangle can be represented by the expression 14x^5 + 63x^2. If the base is 7x^2, write an expression to represent its height.

axie:

My mind broke so like..... @snowflake0531 @carmelle

Vocaloid:

The side length of a cube can be represented by the expression 2x^5. If the side length is doubled, write an expression to represent the new volume of the cube. the original side length is 2x^5. doubling it gives us 2 * 2x^5 or simply 4x^5. from there, the volume is simply the new side length cubed, or (4x^5)^3. simplify.

Vocaloid:

The area of a triangle can be represented by the expression 14x^5 + 63x^2. If the base is 7x^2, write an expression to represent its height. area of a triangle = (1/2)bh so 14x^5 + 63x^2 = (1/2)(7x^2)*h solve for h. first multiply both sides by 2 to eliminate the fraction, then factor out 7x^2 from the left side, then divide both sides by 7x^2. the remaining expression on the right is your height.

Krissy525:

@vocaloid wrote:
The side length of a cube can be represented by the expression 2x^5. If the side length is doubled, write an expression to represent the new volume of the cube. the original side length is 2x^5. doubling it gives us 2 * 2x^5 or simply 4x^5. from there, the volume is simply the new side length cubed, or (4x^5)^3. simplify.
so 64x^15?

Vocaloid:

yup good

Krissy525:

@vocaloid wrote:
yup good
ok

Krissy525:

@vocaloid wrote:
The area of a triangle can be represented by the expression 14x^5 + 63x^2. If the base is 7x^2, write an expression to represent its height. area of a triangle = (1/2)bh so 14x^5 + 63x^2 = (1/2)(7x^2)*h solve for h. first multiply both sides by 2 to eliminate the fraction, then factor out 7x^2 from the left side, then divide both sides by 7x^2. the remaining expression on the right is your height.
what do uu mean by factoring out

Vocaloid:

starting with: 14x^5 + 63x^2 if we factor out 7, we divide whole expression by 7 to get 2x^5 + 9x^5, then we write 7 outside the parentheses 14x^5 + 63x^2 = 7(2x^5 + 9x^2) (think of this like the opposite of applying the distributive property) factoring out x^2 using the same logic 14x^5 + 63x^2 = 7(2x^5 + 9x^2) = 7x^2(2x^3 + 9)

Vocaloid:

from there: area = 7x^2(2x^3 + 9) = (1/2)(7x^2)*h should be straightforward to solve for h from here, cross out the 7x^2 from both sides, multiply both sides by 2

Krissy525:

so it would be h=4x^3+18?

Vocaloid:

yup good

Krissy525:

@vocaloid wrote:
yup good
ok thank you so much